# Learning Priors for Non Rigid SfM from Casual Videos

### v1

@@ 4.3 Architecture
where $L$ is a translation equivariant function, $M_{:,j}\in\mathbb{R}^{N\times d}$ are the columns of $M$ representing all the inputs for a specific tracked point, $F(M)_{:,j}\in\mathbb{R}^{N\times d^{\prime}}$ is the output column and we have $d,d^{\prime}$ input and output feature channels respectively. While valid, this linear layer design might not be the optimal choice. To enhance the model, we incorporate additional nonlinear layers that are well-suited for sequence and set structures data. One particularly effective choice is the transformer layer [46]. Importantly, by leveraging attention mechanisms and temporal positional encodings, the transformer processes sequential input while respecting its ordering and context. Specifically, our layer $F$ is formulated very similarly to Equation (3) and is implemented by replacing the convolutions ($L_{i}$) and summations in Equation (3) with two self-attention mechanisms and suitable temporal positional encoding across the $N$ dimension. We provide more architecture and implementation details in the supplementary.
4.4 Training
Given a dataset of point track representations: $\{M_{1},\dots,M_{L}\}$, our method trains the function $f$ to minimize the loss function over all elements in the dataset:
[…]
Our loss function is defined by:
$\mathcal{L}=\lambda_{\text{Reproject}}\mathcal{L}_{\text{Reproject}}+\lambda_{\text{Rigid}}\mathcal{L}_{\text{Rigid}}+\lambda_{\text{Negative}}\mathcal{L}_{\text{Negative}}+\lambda_{\text{Sparse}}\mathcal{L}_{\text{Sparse}}$ (5)
The first loss is the reprojection error given by the structure and the cameras predicted by $f$:
@@ 4.4 Training
As discussed in Sec. 4.2, the first predicted basis element $B_{1}\in\mathbb{R}^{P\times 3}$ defines a rigid (fixed in time) approximation. Our network also predicts a movement coefficient $\gamma_{j}$ for each world point that defines a zero-mean Cauchy distribution. Given $\gamma_{j}$ and the reprojection error $r_{ij}=\mathcal{R}(B_{1j},R_{i},\mathbf{t}_{i},M_{ij}^{xy})$ of the $j^{th}$ point of $B_{1}$ that is projected by the $i^{th}$ camera, the negative log-likelihood of $r_{ij}$ distributed according to the $\gamma_{j}$-zero-mean Cauchy distribution is proportional to:
@@
Our rigid loss calculates the mean negative log-likelihood over all observed points in all frames:
$\mathcal{L}_{\text{Rigid}}=\frac{1}{\sum_{i=1}^{N}\sum_{j=1}^{P}M_{ij}^{o}}\sum_{i=1}^{N}\sum_{j=1}^{P}M_{ij}^{o}\mathcal{C}(\mathcal{R}(B_{1j},R_{i},\mathbf{t}_{i},M_{ij}^{xy}),\gamma_{j})$ (9)
As in [29] we regularize the points to be in front of the camera by:
@@
We further find it beneficial to regularize the deviation from the rigid approximation $B_{1}$ to be sparse at locations with low $\gamma$ values:
@@
ATE $\downarrow$ RPE Trans $\downarrow$ RPE Rot $\downarrow$ Inference $\downarrow$
(mm) (mm) (deg) Time
DROID-SLAM [43] 5.08 3.60 0.20 10 seconds
ParticleSfM [64] 12.79 6.95 0.51 11 minutes
RCVD[21] 43.95 25.77 2.31 20 minutes
CasualSAM [62] 6.90 3.95 0.22 130 minutes
Ours (cats only) 8.96 3.79 0.23 0.16(+8.6) seconds
Ours (cats only)+BA 4.22 2.86 0.17 0.4(+8.6) seconds
Ours (dogs only) 8.03 3.54 0.23 0.16(+8.6) seconds
Ours (dogs only)+BA 4.19 2.83 0.17 0.4(+8.6) seconds
Ours (dogs+cats) 8.11 3.68 0.24 0.16(+8.6) seconds
Ours (dogs+cats)+BA 4.21 2.86 0.17 0.4(+8.6) seconds
Table 1: Camera Poses Accuracy. Quantitative comparisons between our method to previous methods for camera accuracy and running time. For our method, we compare three versions of (1) training only on cats, (2) training only on dogs, and (3) training on both. We evaluate all methods on an unseen test set of cats and dogs. We further show our accuracies after one round of Bundle Adjustment. For our method’s running time, we also mention the point tracks extraction time in parenthesis (+8.6 seconds) which is performed by [18] as a preprocess.
Abs Rel $\downarrow$ $\delta<1.25\uparrow$ $\delta<1.25^{2}\uparrow$ $\delta<1.25^{3}\uparrow$
Dynamic All Dynamic All Dynamic All Dynamic All
RCVD[21] 0.40 3.6e+7 0.43 0.72 0.75 0.90 0.92 0.96
MiDaS [4] 0.16 6.1e+4 0.78 0.71 0.97 0.88 1.00 0.93
CasualSAM [62] 0.09 0.06 0.93 0.97 0.99 0.99 1.00 1.00
Our (cats only) 0.11 0.08 0.88 0.92 0.99 0.98 1.00 1.00
Our (dogs only) 0.12 0.08 0.85 0.91 0.99 0.99 1.00 1.00
Our (dogs+cats) 0.12 0.08 0.85 0.91 0.98 0.98 1.00 1.00
Table 2: Depth Accuracy We show a comparison to previous methods on the predicted depth for the trajectories compared to their GT depths. We compare three instances of our method, each trained on a different partition of the dataset.
@@
No $\mathcal{L}_{\text{Rigid}}$ 0.77 0.36 0.25 0.46 0.48 0.70 0.68 0.82 1.00 0.86 96.20 29.86 0.99
@@
Table 3: Ablation Study. We evaluate the contribution of different parts from our method by changing one property each time. See details in the text.
Figure 4: Qualitative Results. We present qualitative results in the same format as in Fig. 1. Top. 3 frames from 2 different test video sequences with point tracks marked with corresponding colors. Bottom. A 3D visualization of the outputs of our method, from two time stamps. As the scene is observed from the same viewpoint in each pair, a non-rigid movement can be observed.
@@ 5 Experiments
5.1 Training Details
@@
We trained our network on the cat and dog partitions from the COP3D dataset [38], which contains a diverse set of casual real-world videos of pets. We trained our network from scratch three times to test our generalization capability between semantic categories: once on the cat partition, once on the dog partition, and once on both partitions combined. In total, we used $733$ cat videos and $753$ dog videos for training. We trained our networks for $7000$ and $3500$ epochs for the single-class and multi-class setups respectively. Training our method lasts about one week on a single Tesla V100 GPU with 32GB memory. We used the Adam optimizer [20] with a learning rate of $10^{-4}$. For our non-rigid bases $B_{1},...,B_{K}$ (Sec. 4.2) we used $K=12$. Our method assumes known camera internal parameters which are provided by the dataset and used to normalize the point tracks as a preprocess.
5.2 Evaluation Data
While the COP3D dataset provides cameras that were extracted by COLMAP [37], we note that the non-rigid structure was captured as well in part of their reconstruction which indicates that its reconstruction might not be accurate. Furthermore, the coordinates system units of these reconstructions are unknown. Finally, this dataset does not have any depth map information for evaluating the non-rigid structure.
@@ 5.2 Evaluation Data
For evaluation, we used an RGBD sensor and collected 21 casual videos of dogs and cats, each with 50 frames. The depth maps were used as ground truth for evaluating the reconstructed structure. We extracted the cameras by running COLMAP on the images while masking out the pet areas with dilatated masks provided by [68]. The cameras were scaled to millimeter units using the provided GT depth. We verified that this data includes enough non-rigid motion, by also including several videos that COLMAP failed to reconstruct without the masks. We further verified manually that the camera trajectories look reasonable. Note that this is test data that our network did not see during training and was not used to tune our hyperparameters.
5.3 Results
@@
We evaluated our method on the test data. Qualitative visualizations are presented in Fig. 1 and Fig. 4. We also show a visualization of the movement level values, $\gamma$ in Fig. 5. We compared our method to the state-of-the-art methods and evaluated both, the camera poses and the structure. The camera poses were evaluated using the Absolute Translation Error(ATE), the Relative Translation Error(RTE), and the Relative Rotation Error(RRE) metrics after coordinates system alignment. The results are presented in Tab. 1. We compare three training possibilities of our method of training only on cats, only on dogs, and on both. As can be seen in the table, the results are similar in all 3 cases, even though the dest set contains samples of both dogs and cats. This shows that our network generalizes well to unseen animal classes. Our output camera poses as inferred by the network are already accurate and outperform some of the prior methods. We further show the results of our method after a single and short round of Bundle Adjustment, which makes our method the most accurate one.
Tab. 1 also compares the running times between the methods. It can be seen that our method, even with bundle adjustment is the fastest one when most of the time (8.6 seconds) is used to extract the input point tracks with [18].
To evaluate the structure we compare our method using depth evaluation metrics [13]. The results are presented in Tab. 2, and show that our inferred structure is comparable to the state-of-the-art [62] while taking significantly shorter running times (a few seconds for our method versus more than two hours for [62]). We also note that in contrast to other methods in the table, we are not using any supervision from a depth-from-single-image network.
@@
Figure 5: $\gamma$ Visualization. We show a visualization of the $\gamma$ outputs of our network that are described in Sec. 4.2. In each video sequence, we show the input tracks, where each color visualizes its movement level value, $\gamma$. Purple marks rigid points with low $\gamma$ whereas red marks non-rigid points with high $\gamma$. Note, that our network did not get any direct supervision for these values, but the raw point tracks predictions from [18]. To emphasize our generalization, the $\gamma$ visualizations for cats were produced by the model that was only trained on dogs and vice versa.
5.4 Ablation Study
@@
To evaluate the contribution of our different method parts we run an ablation study which is presented in Tab. 3. In this study, the training was always done on the cats partition from COP3D and evaluated on our test data which contains dogs and cats. More specifically, we ablate on our transformer architecture by taking the architecture suggested by [29] ("Set of Sets") or the DSS architecture [27] ("DSS"). As the table shows our architecture ("Full") achieved significantly better results. We further ignored the $\gamma$ outputs and used regular reprojection error on $B_{1}$ for all points ("No $\gamma$"). We further ablate our sparsity loss ("No $\mathcal{L}_{\text{Sparse}}$") and the rigid loss ("No $\mathcal{L}_{\text{Rigid}}$"). In both cases, the error increased whereas in the later one, the results became much worse.
We further ablate the choice of $K=12$ as the number of non-rigid bases, by trying two extreme numbers of $K=30,K=2$ (we saw no significant differences when we used nearby choices such as $K=11$). As can be seen in the table, when $K=30$ the output is not regularized enough and produces a higher depth error for the dynamic part. For $K=2$ the depth is regularized but the reprojection error ("Rep. (pix.)") gets higher due to over-regularization. Overall the ablation study justifies our design choices ("Full").
5.5 Limitations and Future Work
Our method cannot handle videos with too rapid motion, and in general, is limited by the accuracy of the tracking method [18]. We note that any future improvements with point tracking in terms of accuracy and inference time will directly improve our method as well.
Our method assumes enough motion parallax to constrain the depth values and fails to generate accurate camera poses without it. A future and interesting work would be to try combining depth-from-single-image prior as additional inputs to our network for handling cases with minimal motion parallax and for improving accuracies.
6 Conclusion

### v2

@@ 2.1 Equivariant layers for point track tensors
where $L_{i}$ are linear translation equivariant function (i.e. convolutions), $M_{:,j}\in\mathbb{R}^{N\times d}$ are the columns of $M$ representing all the inputs for a specific tracked point, $F(M)_{:,j}\in\mathbb{R}^{N\times d^{\prime}}$ is the output column and $d,d^{\prime}$ are the input and output feature channels respectively. To construct a neural network, these layers can be interleaved with pointwise nonlinearities, similar to basic convolutional neural networks.
Implementation via transformers and positional encoding.
While the linear layer design is reasonable, it may not be the optimal choice. To enhance the model, we design a new layer whose structure follows Equation (1), but incorporates nonlinear layers in the form of transformers [47]. Specifically, our layer $F$ is formulated similarly to Equation (1), but instead of convolutions ($L_{i}$) and summations, it utilizes two self-attention mechanisms and suitable temporal positional encoding across the $N$ dimension. Formally, our basic layer $F:\mathbb{R}^{N\times P\times d}\rightarrow\mathbb{R}^{N\times P\times d^{\prime}}$ is computed via four steps, which are described below:
$\mathbf{\bar{q}}_{ij}=\bar{W}^{Q}M_{ij},~{}\mathbf{\bar{k}}_{ij}=\bar{W}^{K}M_{ij},~{}\mathbf{\bar{v}}_{ij}=\bar{W}^{V}M_{ij}\Rightarrow\bar{M}_{ij}=\sum_{i^{\prime}=1}^{N}\frac{\exp(\mathbf{\bar{q}}_{ij}\cdot\mathbf{\bar{k}}_{i^{\prime}j})}{\sum_{l=1}^{N}\exp(\mathbf{\bar{q}}_{ij}\cdot\mathbf{\bar{k}}_{lj})}\mathbf{\bar{v}}_{i^{\prime}j}$ (2)
$\mathbf{q}_{ij}=W^{Q}\bar{M}_{ij},~{}\mathbf{k}_{ij}=W^{K}\bar{M}_{ij},~{}\mathbf{v}_{ij}=W^{V}\bar{M}_{ij}\Rightarrow F(M)_{ij}=\sum_{j^{\prime}=1}^{P}\frac{\exp(\mathbf{q}_{ij}\cdot\mathbf{k}_{ij^{\prime}})}{\sum_{l=1}^{P}\exp(\mathbf{q}_{ij}\cdot\mathbf{k}_{il})}\mathbf{v}_{ij^{\prime}}$ (3)
Here, $M_{i,j}\in\mathbb{R}^{d}$ are the features associated with the $j$-th point in the $i$-th frame. The attention mechanism defined in the first equation above (2) is augmented with standard temporal positional encoding in the first layer and replaces the translation equivariant function $L_{i}$ applied to the columns of $M$ (Eq.(1)). The attention in the second equation (3) implements the set aggregation (summation) (also in Eq.(1)) applied to the rows of $M$. As commonly done, we use transformers with 16 attention heads [47].
2.2 Constraining 3D motion and camera poses via low-rank assumption
Given our 2D tracks, we aim to characterize the motion of the points by decomposing them into the global camera motion and the 3D motion of objects in the scene. The 2D motion of static scene points provides useful constraints for estimating the camera motion. However, as previously mentioned, predicting camera and dynamic 3D motion solely from 2D motion is an ill-posed problem without additional constraints [3]. We tackle this challenge by adding two mechanisms to our architecture: (1) low-rank movement assumption; and (2) specific modeling of the static scene for camera estimation.
Low-rank movement assumption.
First, motivated by classical orthographic Non-Rigid Structure from Motion [6], we constrain the output points to be formulated by a linear combination of input-specific basis elements. Specifically, given the input 2D point tracks, $M\in\mathbb{R}^{N\times P\times 3}$, our network predicts $K$ point clouds: $B_{1}(M),\dots,B_{K}(M)\in\mathbb{R}^{P\times 3}$ and $N(K-1)$ linear coefficients, $\{c_{1k}(M)\}_{k=2}^{K},\dots\{c_{Nk(M)}\}_{k=2}^{K}$ such that $X_{i}(M)=B_{1}(M)+\sum_{k=2}^{K}c_{ik}(M)B_{k}(M),$ where $X_{i}(M)\in\mathbb{R}^{P\times 3}$ is the 3D point cloud at frame $i$. The point clouds and coefficients are computed by taking the output of the last equivariant layer as defined in the previous section and applying invariant aggregations on the respective dimension resulting in equivariant and invariant outputs. See more details in the appendix. We note that we deliberately chose the coefficient of $B_{1}(M)$ to be the constant $1$, the reason is explained in the next paragraph.
Specific modeling of the static scene for camera estimation.
Frequently, casual video data of dynamic scenes contains many static regions, which can be used to determine camera poses [61]. We leverage this observation by treating the first basis element $B_{1}(M)\in\mathbb{R}^{P\times 3}$ as a static approximation for all scene points and encourage $B_{1}(M)$ as well as the output camera poses to explain the 2D observations according to this approximation using a "static" reprojection loss ($\mathcal{L}_{\text{Static}}$, defined in the next section). We note, however, that a static point cloud is not likely to produce low reprojection errors for the non-static components, thus the reprojection error necessitates robustness to substantial errors from the non-static elements. To address this, our network predicts (equivariantly) $P$ motion level values $\gamma_{1}(M),\dots,\gamma_{P}(M)\in\mathbb{R}_{+}$, one for each point in our dynamic point cloud, which we use to weight the reprojection errors from $B_{1}(M)$. The main idea is to give less weight to non-static points so that the static projection loss can disregard them. Specifically, inspired by [59], each $\gamma_{i}(M)$ defines a Cauchy distribution that models the reprojection errors for its associated world point, such that a world point with higher $\gamma$ is expected to produce a wider error distribution. Empirically, as noted by [59], the Cauchy distribution tends to be more robust for modeling reprojection error uncertainties compared to Gaussian noise modeling [19]. Then, $\mathcal{L}_{\text{Static}}$, minimizes the negative log-likelihood under this assumption. See details in Sec. 2.3.
2.3 Training and losses
Model outputs. Given the input 2D point tracks $M\in\mathbb{R}^{N\times P\times 3}$, our network produces outputs as a function of $M$: linear bases and coefficients $B_{1}(M),\dots,B_{K}(M)\in\mathbb{R}^{P\times 3},\{c_{1k}(M)\}_{k=2}^{K},\dots,\{c_{Nk}(M)\}_{k=2}^{K}\in\mathbb{R}$ which define a dynamic point cloud $X_{1}(M),\cdots,X_{N}(M)\in\mathbb{R}^{P\times 3}$,$\gamma_{1}(M),\dots,\gamma_{P}(M)\in\mathbb{R}_{+}$ movement level values, and $(R_{1}(M),\mathbf{t}_{1}(M)),\dots,(R_{N}(M),\mathbf{t}_{N}(M))\in SO(3)\times\mathbb{R}^{3}$ camera poses.
We use these network outputs to define an self-supervised loss function with respect to the current network weights and $M$ which is defined by:
$\mathcal{L}=\lambda_{\text{Reproject}}\mathcal{L}_{\text{Reproject}}+\lambda_{\text{Static}}\mathcal{L}_{\text{Static}}+\lambda_{\text{Negative}}\mathcal{L}_{\text{Negative}}+\lambda_{\text{Sparse}}\mathcal{L}_{\text{Sparse}}$ (4)
Reprojection loss.
The reprojection loss encourages the consistency between the output 3D point clouds and camera poses, to the input 2D observations:
@@ 2.3 Training and losses
Static loss.
As discussed in Sec. 2.2, to better constrain the camera poses, the first predicted basis element $B_{1}(M)\in\mathbb{R}^{P\times 3}$ defines a static (fixed in time) point cloud approximation. Our network also predicts a movement coefficient $\gamma_{j}(M)$ for each world point that defines a zero-mean Cauchy distribution. Given $\gamma_{j}$ and the reprojection error $r_{ij}=\mathcal{R}(B_{1j},R_{i},\mathbf{t}_{i},M_{ij}^{xy})$ of the $j^{th}$ point of $B_{1}$ that is projected by the $i^{th}$ camera, the negative log-likelihood of $r_{ij}$ distributed according to the $\gamma_{j}$-zero-mean Cauchy distribution is proportional to:
@@
Note, that this loss reduces the contribution of the reprojection errors for points with high $\gamma$, but also encourages $\gamma$ to be small, i.e. encouraging the static point cloud to represent the dynamic scene when possible. Our static loss is the mean negative log-likelihood over all observed points in all frames:
$\mathcal{L}_{\text{Static}}=\frac{1}{\sum_{i=1}^{N}\sum_{j=1}^{P}M_{ij}^{o}}\sum_{i=1}^{N}\sum_{j=1}^{P}M_{ij}^{o}\mathcal{C}(\mathcal{R}(B_{1j},R_{i},\mathbf{t}_{i},M_{ij}^{xy}),\gamma_{j})$ (8)
Regularization losses.
As in [29] we regularize the observed points to be in front of the camera by:
@@
We further find it beneficial to regularize the deviation from the static approximation $B_{1}$ to be sparse for static points, i.e. points with low $\gamma$ values:
@@
Table 1: Pet evaluation. Top: Baseline method results for structure or camera estimation (or both). Bottom: Our results with several configurations. (C),(D), or (CD) respectively indicate the object categories in the training set: cats, dogs, or both. BA and FT respectively indicate a post-processing of Bundle Adjustment or fine-tuning.
Abs Rel $\downarrow$ $\delta<1.25\uparrow$ $\delta<1.25^{2}\uparrow$ $\delta<1.25^{3}\uparrow$ ATE $\downarrow$ RPE Trans $\downarrow$ RPE Rot $\downarrow$ Time
Dyn. All Dyn. All Dyn. All Dyn. All (mm) (mm) (deg) (min)
D-SLAM [44] - - - - - - - - 5.08 3.60 0.20 0.16
ParticleSFM [61] - - - - - - - - 12.79 6.95 0.51 11.00
RCVD [22] 0.40 3.6E+07 0.43 0.72 0.75 0.90 0.92 0.96 43.95 25.77 2.31 20.00
CasualSAM [59] 0.09 0.06 0.93 0.97 0.99 0.99 1.00 1.00 6.90 3.95 0.22 1.3E+02
MiDaS [4] 0.16 6.2E+04 0.78 0.71 0.97 0.88 1.00 0.93 - - - 0.15
Ours (C) 0.11 0.08 0.88 0.92 0.99 0.98 1.00 1.00 8.96 3.79 0.23 0.15
Ours (C)+BA 0.11 0.08 0.88 0.92 0.99 0.98 1.00 1.00 4.22 2.86 0.17 0.15
Ours (C)+FT 0.09 0.06 0.90 0.96 1.00 0.99 1.00 1.00 4.00 2.74 0.16 4.86
Ours (D) 0.12 0.08 0.85 0.91 0.99 0.99 1.00 1.00 8.03 3.54 0.23 0.15
Ours (D)+BA 0.12 0.08 0.85 0.91 0.99 0.99 1.00 1.00 4.19 2.83 0.17 0.15
Ours (D)+FT 0.09 0.06 0.88 0.96 1.00 0.99 1.00 1.00 3.98 2.74 0.16 4.86
Ours (CD) 0.12 0.08 0.85 0.91 0.98 0.98 1.00 1.00 8.11 3.68 0.24 0.15
Ours (CD)+BA 0.12 0.08 0.85 0.91 0.98 0.98 1.00 1.00 4.21 2.86 0.17 0.15
Ours (CD)+FT 0.09 0.06 0.90 0.96 1.00 0.99 1.00 1.00 3.98 2.74 0.16 4.86
Table 2: Out-of-training-domain evaluation . Evaluation metrics on monocular videos from [58]. The table has the same structure as Tab. 1.
Abs Rel $\downarrow$ $\delta<1.25\uparrow$ $\delta<1.25^{2}\uparrow$ $\delta<1.25^{3}\uparrow$ ATE $\downarrow$ RPE Trans $\downarrow$ RPE Rot $\downarrow$ Time
Dyn. All Dyn. All Dyn. All Dyn. All (mm) (mm) (deg) (min)
D-SLAM [44] - - - - - - - - 7.96 10.91 0.07 0.18
ParticleSFM [61] - - - - - - - - 26.66 23.83 0.20 2.13
RCVD [22] 0.19 2.6E+05 0.69 0.75 0.95 0.95 0.96 0.98 1.6E+02 3.2E+02 3.43 7.00
CasualSAM [59] 0.05 0.03 0.95 0.99 0.99 1.00 1.00 1.00 7.81 10.09 0.06 22.00
MiDaS [4] 2.8E+04 2.7E+05 0.59 0.58 0.73 0.72 0.83 0.80 - - - 0.02
Ours (C) 0.08 0.06 0.89 0.95 0.99 0.99 0.99 1.00 32.06 47.99 0.45 0.04
Ours (C)+BA 0.08 0.06 0.89 0.95 0.99 0.99 0.99 1.00 8.67 12.36 0.08 0.04
Ours (C)+FT 0.07 0.03 0.94 0.98 0.99 1.00 1.00 1.00 7.98 11.64 0.08 0.59
Ours (D) 0.08 0.07 0.92 0.93 0.99 0.98 0.99 1.00 33.77 51.64 0.61 0.04
Ours (D)+BA 0.08 0.07 0.92 0.93 0.99 0.98 0.99 1.00 8.40 12.06 0.08 0.04
Ours (D)+FT 0.05 0.03 0.97 0.99 0.99 1.00 0.99 1.00 8.15 11.88 0.09 0.59
Ours (CD) 0.10 0.08 0.93 0.94 0.99 0.99 1.00 1.00 36.17 53.94 0.67 0.04
Ours (CD)+BA 0.10 0.08 0.93 0.94 0.99 0.99 1.00 1.00 8.62 12.49 0.08 0.04
Ours (CD)+FT 0.06 0.03 0.97 0.99 0.99 1.00 0.99 1.00 8.04 11.84 0.09 0.59
Table 3: Ablation study. The contribution of different parts from our method. See details in the text.
@@
No $\mathcal{L}_{\text{Static}}$ 0.77 0.36 0.25 0.46 0.48 0.70 0.68 0.82 1.00 0.86 96.20 29.86 0.99
@@ 3 Experiments
In this section, we conduct experiments to verify our proposed network’s performance on real-world casual videos. We began by training the network on specific domains and then evaluated its accuracy and running time on unseen videos from both, training and unseen domains.
@@
Training procedure. We trained our network on the cat and dog partitions from the COP3D dataset [40], which contains a diverse set of casual real-world videos of pets. Our networks were trained from scratch three times to test our generalization capability between semantic categories: once on the cat partition, once on the dog partition, and once on both partitions combined. Training technical details are provided in the appendix. We use $K=12$ bases in all our experiments (Sec. 2.2).
@@
Evaluation data. To assess our framework’s performance on pet videos, we curated a new dataset333 While the COP3D dataset provides cameras that were extracted by COLMAP [38], we note that this evaluation data is insufficient in our case. This is because the dynamic structure was captured as well in part of their reconstruction which indicates that its reconstruction might not be accurate. Furthermore, the coordinates system units of these reconstructions are unknown. Finally, this dataset does not have any depth map information for evaluating the dynamic structure. consisting of 21 casual videos of dogs and cats, each video containing 50 frames. These videos were captured using an RGBD (RGB-Depth) sensor. The depth maps were used as ground truth for evaluating the reconstructed structure. We extracted the cameras by running COLMAP on the images while masking out the pet areas with dilatated masks provided by [65]. The cameras were scaled to millimeter units using the provided GT depth. Note that our network did not see this test data during training and it was not used to tune our hyperparameters.
Additionally, to evaluate our method on out-of-domain evaluation data, we used the Nvidia Dynamic Scenes Dataset [58]. Specifically, while our network was trained on pet videos, this dataset contains other dynamic object types, e.g. human, balloon, truck, and umbrella, with a different camera motion type. The dataset contains 8 dynamic scenes which are captured by 12 synchronized cameras, enabling accurate depth estimation which is treated as GT for evaluating monocular depth estimation. The ground truth camera poses were calculated by [38] with the synchronized multiview camera rig and the ground truth dynamic masks. Similarly to [24] we simulated 8 monocular dynamic video sequences using the camera rig, each with 24 frames, and used them for evaluation.
Evaluation results.
@@
Qualitative visualizations are presented in Fig. 3.444The reader is encouraged to watch the supplementary videos for better 4D perception. We also show a visualization of the movement level values, $\gamma$ in Fig. 4. For comparisons, we chose state-of-the-art methods that as our method, can be applied to raw casual videos that were captured by standard pinhole camera models and do not need any static or dynamic segmentation. We evaluated both, the camera poses and the structure accuracies. Comparison results for the pet-test-set and out-of-domain dataset are presented in Tables 1 and 2 respectively. The camera poses are evaluated compared to the GT, using the Absolute Translation Error(ATE), the Relative Translation Error(RTE), and the Relative Rotation Error(RRE) metrics after coordinates system alignment. We compare three training configurations of our method of training only on cats, only on dogs, and on both. As can be seen in the tables, the results are similar in all 3 cases. Our output camera poses as inferred by the network are already accurate and outperform some of the prior methods. We further show the results of our method after a single and short round of Bundle Adjustment, which makes our method better than all baselines on the pet sequences, and comparable on the out-of-domain cases.
Importantly, Tables 1 and 2 also compare the method’s running times. It can be seen that our method, even with bundle adjustment, is the fastest camera prediction method. Tables 1 and 2 also show structure evaluation with the depth evaluation metrics [11] on the sampled point tracks. They demonstrate that our inferred structure is almost comparable to the state-of-the-art [59] while taking significantly shorter running times (a few seconds for our method versus more than two hours for [59] on pet videos). Short (0.6-5 minutes), per-sequence fine-tuning makes our method’s accuracy comparable to [59], see appendix for details. In terms of running time, our method is a bit slower than MiDaS [4], which only provides depth maps without cameras, but achieves much better results. We note that in contrast to the other methods that predict the dynamic depth, ours does not use any depth-from-single-image prior.
Figure 3: Qualitative Results. Top. Frames from 2 different test video sequences with point tracks marked with corresponding colors. Bottom. A 3D visualization of our method’s outputs, from two time stamps. The camera trajectory is present as gray frustums, whereas the current camera is marked in red. The reconstructed 3D scene points are presented in corresponding colors to the input tracks on the top. The scene is observed from the same viewpoint, enabling the visualization of the dynamic reconstructed structure.
@@
Figure 4: $\gamma$ Visualization. We show a visualization of the $\gamma$ outputs of our network that are described in Sec. 2.2. In each video sequence, we show the input tracks, where each color visualizes its movement level value, $\gamma$. Purple marks static points with low $\gamma$ whereas red marks dynamic points with high $\gamma$. Note, that our network did not get any direct supervision for these values, but only the raw point tracks predictions from [18]. The $\gamma$ visualizations for cats were produced by the model that was only trained on dogs and vice versa. We note that our model generalizes well to out-of-domain (non-pet) cases as well.
Ablation study
@@
To evaluate the contribution of our different method parts we run an ablation study which is presented in Tab. 3. In this study, the training was always done on the cat partition from COP3D and evaluated on our test data which contains dogs and cats. First, we performed an ablation study on our transformer architecture by taking the architecture suggested by [29] ("Set of Sets") or the DSS architecture that uses only linear layers [27] ("DSS"). As the table shows our architecture ("Full") achieved significantly better results. To test the losses in our framework, we also evaluated the following: (1) ignoring the $\gamma$ outputs and using regular reprojection error on $B_{1}$ for all points ("No $\gamma$"); (2) removing our sparsity loss ("No $\mathcal{L}_{\text{Sparse}}$"); and (3) removing the static loss ("No $\mathcal{L}_{\text{Static}}$"). In all cases, the error increased whereas in the later one, the results became much worse. We further ablate the choice of $K=12$ as the number of linear bases, by trying 2 extreme numbers of $K=30,K=2$ (we saw no significant differences when we used nearby choices such as $K=11$). As can be seen in the table, when $K=30$ the output is not regularized enough and produces a higher depth error for the dynamic part. For $K=2$ the depth is regularized but the reprojection error ("Rep. (pix.)") gets higher due to over-regularization. Overall, this study justifies our design choices ("Full").
4 Related Work
Simultaneous Localization and Mapping (SLAM) and Structure from Motion (SfM)
SfM pipelines seek to recover static 3D structure and camera poses from unordered images.[45, 41, 38, 1, 51]. Learning-free pipelines [38] are effective but require repeated applications of Bundle Adjustment (BA) [46]. [29, 8] presented a method for learning prior from a dataset of multiview image sets, to accelerate SfM pipelines by using equivariant deep networks. Monocular Simultaneous Localization and Mapping (SLAM) methods [30, 31, 12, 57, 5, 50, 62, 63, 42] extract camera poses from video sequences while defining a scene map with keyframes. These methods assume static scenes, fail to produce the cameras in scenes with large portions of dynamic motion, and cannot reproduce dynamic parts of the scene. DROID-SLAM [44] used synthetic data with ground truth 3D supervision for learning to predict camera poses via deep-based BA on keyframes while excluding dynamic objects. ParticleSfM [61] filters out 2D dynamic content for reproducing the cameras in dynamic scenes, using its pre-trained motion prediction network. Both, [44, 61] do not infer the dynamic 3D structure.
Orthographic Non-Rigid SfM (NRSfM)
Bregler et al. [6] introduced a factorization method for computing a non-rigid structure and rotation matrices from a point track matrix, by assuming a low dimensional basis model. While follow-up papers improved accuracies with different regularizations [23, 9, 33, 16] or neural representations [32, 21, 39], the orthographic camera model assumption is in general not valid for casual videos. Furthermore, these methods often assume background subtraction as a preprocessing. Even though a follow-up work proposed factorization solutions for pinhole cameras [14], its sensitivity to noise [17], makes it impractical for casual videos.
Test-time optimization for dynamic scenes
Recent methods [26, 22, 59, 60] finetuned the monocular depth estimation from a pre-trained model [37, 36] using optical flow constraints [43], for obtaining consistent dense depth maps for a monocular video. [59] further optimized motion maps for handling scenes with large dynamic motion. [52, 13] use depth from single image estimations, to improve novel view synthesis in dynamic scenes. [25] further optimizes for the unknown camera poses together with the dynamic radiance field optimization. [34, 35] model a single deformable surface from a monocular video by applying isometric constraints. LASR [54], ViSER [55] and BANMo [56] optimize for a dynamic surface by assuming rigid bones and linear blend skinning weights. However, all the above-mentioned methods require per-scene optimization, resulting in slow inference. Recently, [40] presented the Common Pets in 3D (COP3D) dataset that contains casual, in-the-wild videos of pets, and used it to learn priors for novel view synthesis in dynamic scenes.
Point tracking
There has been a recent advance in 2D point tracking by learning [18, 10], or optimization [49] techniques. Concurrently, [53] presented a method for jointly performing 2D tracking and 3D lifting, by learning to track with depth supervision while applying an as-rigid-as-possible loss. However, their method cannot predict camera poses or identify static parts directly.
5 Conclusions and limitations